OpenStudy (anonymous):

tell wheather the ordered pair (-3,2) is a solution of the system (2x+y=-4 3x-y+-11)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

i am assuming the second equation is 3x-y=11 and not 3x-y+-11 ?

OpenStudy (anonymous):

3x-y=-11

OpenStudy (anonymous):

okay. do you know how to solve a system of equations?

OpenStudy (anonymous):

not really im a blonde at math !

OpenStudy (amistre64):

(x=-3,y=2) plug them in and see if they match :)

OpenStudy (anonymous):

2x+y = -4 ----------------------------- equation 1 3x-y = -11 ---------------------------- equation 2 add the two equations 2x+y + (3x -y) = -4 + (-11) 5x = -15 x = -3 substitute in equation 1 2(-3) + y = -4 -6+y = -4 y = 2 so x = -3, y = 2 or the ordered pair (-3,2) is the solution for the system of equations.